OpenStudy (anonymous):

What is the geometric mean of 3 and 7?

OpenStudy (anonymous):

sqrt of multiple

OpenStudy (anonymous):

^^ I Agree With Estudier

OpenStudy (anonymous):

so whats the answer then

OpenStudy (anonymous):

He says its the square root of the multiple right so would it be the square root of 21?

OpenStudy (anonymous):

So, GM of a,b,c..... is sqrt(a*b*c......)

OpenStudy (anonymous):

I might be wrong but I thought it was the nth root if there are n numbers. For the case of two numbers, such as 3 and 7, square root would do. But for the GM of a, b,c... wouldn't it be nth root of (a*b*c*....etc)?

OpenStudy (anonymous):

Do You Understand

OpenStudy (anonymous):

Yes, ur right, not sqrt, kth root or whatever....sorry

OpenStudy (anonymous):

No I do not. I am terrible at Geometry!

OpenStudy (anonymous):

Okay Hold On

OpenStudy (anonymous):

kk

hartnn (hartnn):

Geometric Mean of a and b is \(\sqrt{ab}\) so geometric mean of 3 and 7 will be \(\huge \sqrt{3*7}=\sqrt{21}=?\)

OpenStudy (anonymous):

I have to find the square root of 21.

OpenStudy (anonymous):

its 4.5???

OpenStudy (anonymous):

4.582575694955840006588047193728

hartnn (hartnn):

if u want numerical answer then yes, find the square root, else if the answer can be in radical form, then leave it as \(\sqrt{21}\)

OpenStudy (anonymous):

These are the multiple choice answers √21 2√6 2√7 21

hartnn (hartnn):

so its 1st one, isn't it?

OpenStudy (anonymous):

The Answer Is The Second One ,

OpenStudy (anonymous):

why the second? we want sqrt(21).. which is the first choice

OpenStudy (anonymous):

how do you know it's the second one?

OpenStudy (anonymous):

Im Sorry It Is The 1st One Im Sorry ,

OpenStudy (anonymous):

thanks everybody for you help :)

hartnn (hartnn):

no problem, 1st one it is.

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

what about for 18 and 8

OpenStudy (anonymous):

the geometric mean for 18 and 8... I would first multiply then square root

OpenStudy (anonymous):

what do you think about that one? if you follow the same method, what should you get?

OpenStudy (anonymous):

yes

hartnn (hartnn):

\(\huge \sqrt {18*8}=?\)

OpenStudy (anonymous):

12

OpenStudy (anonymous):

yes good

OpenStudy (anonymous):

i mean 144 but then square root it, it would equal 12

hartnn (hartnn):

12 it is :)

OpenStudy (anonymous):

Alright thanks!

OpenStudy (anonymous):

just make sure, that if there are more than two numbers, you take the appropriate root. :)

OpenStudy (anonymous):

Can I ask you guys another question

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

What are the values of a and b? a = 20, b = 2√101 a = 20, b = 20√101 a = 30, b = 2√101 a = 400, b = 404

OpenStudy (anonymous):

wait I need to put a picture

OpenStudy (anonymous):